# SVEMnet: An R package for Self-Validated Elastic-Net Ensembles and Multi-Response Optimization in Small-Sample Mixture–Process Experiments

### v1

@@ Abstract
SVEMnet is an R package for fitting Self-Validated Ensemble Models (SVEM) with elastic-net base learners and for performing multi-response optimization in small-sample mixture–process design-of-experiments (DOE) studies with numeric, categorical, and mixture factors. SVEMnet wraps elastic-net and relaxed elastic-net models for Gaussian and binomial responses from glmnet in a fractional random-weight (FRW) resampling scheme with anti-correlated train/validation weights; penalties are selected by validation-weighted AIC- and BIC-type criteria, and predictions are averaged across replicates to stabilize fits near the interpolation boundary. In addition to the core SVEM engine, the package provides deterministic high-order formula expansion, a permutation-based whole-model test heuristic, and a mixture-constrained random-search optimizer that combines Derringer–Suich desirability functions, bootstrap-based uncertainty summaries, and optional mean-level specification-limit probabilities to generate scored candidate tables and diverse exploitation and exploration medoids for sequential fit–score–run–refit workflows. A simulated lipid nanoparticle (LNP) formulation study illustrates these tools in a small-sample mixture–process DOE setting, and simulation experiments based on sparse quadratic response surfaces benchmark SVEMnet against repeated cross-validated elastic-net baselines.
@@ 1 Introduction
This paper introduces SVEMnet, an R package that implements SVEM with elastic-net and optional relaxed elastic-net base learners [8, 9, 10, 11]. In a relaxed fit, glmnet computes a penalized solution path indexed by the penalty parameter $\lambda$, identifies the active set at each $\lambda$, and then refits an unpenalized least‑squares (or GLM) model on that active set, interpolating between penalized and unpenalized coefficients via a relaxation parameter $\gamma\in[0,1]$. In the sparse Gaussian linear model studied by Meinshausen [10], the relaxed lasso reduces the shrinkage of non-zero coefficients and can achieve lower prediction error than the standard lasso. Simulations in Section 4 indicate that combining relaxed base learners with the SVEM FRW ensemble can further reduce out-of-sample prediction error for Gaussian responses, plausibly because the ensemble averages over the additional variance introduced by relaxation.
@@
Self‑Validated Ensemble Models have already seen practical use in industrial DOE and chemometric settings, including mixture–process workflows for lipid nanoparticle (LNP) formulations and other pharmaceutical and analytical applications [17, 18, 6, 19, 20, 21, 22, 23, 24]. These applications motivate an implementation tailored to chemometric DOE. In particular, SVEMnet is designed to automate many of the steps in the SVEM-based mixture–process workflow for LNP formulation optimization described by Karl et al. [17]. SVEMnet supplies an open-source SVEM implementation over glmnet with validation-weighted AIC/BIC selectors, deterministic high-order full model expansions, a permutation-based whole-model test heuristic, and a mixture-constrained random-search scoring tool for multi-response desirability and candidate generation.
@@ 2.2 SVEM background
following the FRW scheme for univariate analysis introduced by Xu et al. [3], which underpins the original SVEM extensions to penalized regression by Gotwalt and Ramsey [4] and Lemkus et al. [2]. For each FRW replicate we then rescale both the training and validation weights to have mean one, $w_{i}\leftarrow w_{i}/\bar{w}$ for $w\in\{w^{\mathrm{train}},w^{\mathrm{valid}}\}$, matching glmnet’s internal normalization of observation weights.
Given $(w^{\mathrm{train}},w^{\mathrm{valid}})$ for replicate $b$, SVEMnet fits an elastic-net or relaxed elastic-net base learner on $(X,y)$ under $w^{\mathrm{train}}$, evaluates a collection of path points by a validation-weighted information criterion computed with $w^{\mathrm{valid}}$ (Section 2.4), and stores predictions from the selected path point. Overall SVEM predictions are obtained by averaging these member predictions across $b=1,\dots,B$. Lemkus et al. [2] tuned penalties by validation-weighted SSE (wSSE); SVEMnet adopts the same FRW-and-ensemble structure but standardizes on validation-weighted AIC/BIC analogs—wAIC default for Gaussian responses; wBIC default for binomial responses.
@@ 2.3 Implementation note: a SVEM wrapper over glmnet
For new data, the design matrix is rebuilt to match training, each replicate’s coefficients are applied, and the resulting member predictions are averaged—on the response scale for Gaussian models and, by default, on the probability scale for binomial models.
In this sense, penalized optimization is delegated to glmnet; SVEMnet contributes the FRW-based self-validation and ensembling, the validation-weighted AIC/BIC analogs with effective-validation-size guardrails, and the DOE-oriented expansion, testing, and candidate-generation tools used in the workflows of Section 3.
@@ 2.5.2 Scoring and response weighting
User-specified goal weights $\tilde{w}_{r}$ are first normalized to $\bar{w}_{r}$ with $\sum_{r}\bar{w}_{r}=1$. When reweight_by_wmt = FALSE, we set $w_{r}=\bar{w}_{r}$ and report a single column score based on $S(x)$. When reweight_by_wmt = TRUE for Gaussian responses, we further reweight the normalized user weights by a monotone function of the permutation whole-model $p$-values from [5] (for example $\log_{10}(p_{r})$), renormalize to obtain WMT-adjusted weights $w_{r}^{\mathrm{WMT}}$, and compute a second geometric-mean score wmt_score by using $w_{r}=w_{r}^{\mathrm{WMT}}$ in $S(x)$. These WMT-adjusted weights are diagnostic and do not alter the underlying SVEM fits.
@@ 5 Ionizable_Lipid_Type + N_P_ratio + flow_rate,
9 polynomial_order = 3, # up to cubic terms I(X^2), I(X^3)
@@ 3.1 Whole-model reweighting and permutation tests
To assess which responses show the strongest relationship with the study factors, we apply the permutation-based whole-model test heuristic (WMT) of Karl [5] to each SVEM model. WMT-based multipliers can optionally reweight the user-specified response weights, prioritizing responses with stronger predictive signal and down-weighting weaker ones. If utilized, WMT-reweighted candidate sets should be secondary to those formed without the WMT reweighting, since the heuristic may have limited power in some situations [5]. Because WMT is permutation-based with finite simulation budgets, these $p$-values are approximate.
@@ 5 PDI = form_pdi),
Table 1: Permutation whole-model test results for the three responses (SVEMnet).
@@
Size $1\text{\times}{10}^{-16}$
Potency $3\text{\times}{10}^{-7}$
PDI 0.78
Notes. Values are permutation-based and subject to Monte Carlo error.
@@ 3 Size = list(upper = 100),
The random-search scoring step then evaluates a large number of feasible settings (here 25,00025{,}000). For each candidate, svem_score_random() evaluates the SVEM models, computes per-response Derringer–Suich desirabilities, aggregates them into a single multi-response score via a weighted geometric mean, constructs a scalar uncertainty measure $U(x)$ from bootstrap percentile intervals, and, when specifications are supplied, computes mean-level probabilities that each candidate’s modeled mean lies inside the user limits. Supplying wmt = wmt_out adds a WMT-adjusted score column wmt_score in addition to the user-weighted score:
@@ 3.3 Computational environment and runtime
All timings were obtained on a Windows 11 x64 desktop with an Intel Core Ultra 7 265 CPU (20 cores, 2.4 GHz) and 64 GB RAM, running R 4.5.2 with SVEMnet 3.1.4 and glmnet 4.1-10 and parallel execution enabled for the WMT. Users should note that on Windows, the parallelization of the WMT will trigger a one-time Windows Firewall prompt for R when worker processes are created; this is due to local socket connections between R processes and does not involve external network access.
@@ 4.1 Design and metrics
Each replicate uses a sparse quadratic response surface in four continuous factors $X_{1}{:}X_{4}\in[-1,1]$ and one three-level factor $X_{5}\in{\mathrm{L1},\mathrm{L2},\mathrm{L3}}$ (sum-to-zero coding). Training data are Latin hypercube samples for $X_{1}{:}X_{4}$ with balanced $X_{5}$ levels [31], with $n_{\text{total}}\in\{15,20,\dots,50\}$. As noted in the introduction, we use Latin hypercubes here for simulation convenience; in practice, more sophisticated optimal or space-filling designs are typically preferable when constructing the initial experimental design in applied settings [39, 27, 29, 30, 28].
@@
To reduce Monte Carlo error, the summary curves in Figures 3, 4, and 5 average over 900 independent simulation replicates for each combination of $n_{\text{total}}$, target $R^{2}$, model order, and fitting method (Setting). In these overview plots we fix a common set of tuning defaults. For Gaussian responses, cv.glmnet uses standard (non-relaxed) elastic-net paths (relax = FALSE) and SVEMnet uses relaxed base learners (relax = TRUE); post-hoc debiasing is disabled for all methods and the elastic-net mixing grid is held at $\alpha\in\{0.5,1\}$. For binomial responses, all methods use non-relaxed paths (relax = FALSE) with the same default $\alpha$ grid. These settings are used for all model orders shown in the figures.
@@ 4.2 Gaussian results
Figure 3: Effect of expansion order on mean log-NRMSE (lower is better) on a noiseless holdout. Panels: target $R^{2}\in\{0.5,0.9\}$. Columns: model order used in fitting. Curves compare repeated cv.glmnet baselines (standard, non-relaxed paths) to SVEMnet with relaxed elastic-net base learners using the default mixing grid $\alpha\in\{0.5,1\}$; no debiasing is applied. Underspecification (Order 1) consistently harms accuracy; modest overspecification (Order 3) degrades far less than underspecification, with Order 2 best as expected.
Figure 4: Mean median-selected size $k_{\lambda}$ versus $n_{\text{total}}$ for Gaussian simulations, stratified by target $R^{2}$ and model order. Colors denote selector (cv.glmnet, wAIC, wSSE, wBIC); line type denotes $\alpha\in\{0.5,1\}$. Simulation settings match Figure 3 (non-relaxed cv.glmnet, relaxed SVEMnet, no debiasing). wSSE exhibits a localized spike in $k_{\lambda}$ near $n=p_{\mathrm{full}}$.
@@ 4.3 Binomial results
Across $n_{\text{total}}\in\{15,\dots,50\}$ and $R^{2}\in\{0.5,0.9\}$, non-relaxed fits dominate their relaxed counterparts on the holdout log-loss scale. For the correctly specified quadratic expansion (Order 2), SVEMnet with wBIC and the non-relaxed repeated-cv.glmnet baselines attain essentially indistinguishable performance: Tukey–Kramer comparisons based on the mixed-effects model place these configurations in the same best-performing group (Table LABEL:tab:tukey-binomial). However, Figure 5 suggests that SVEMnet with wBIC may outperform cv.glmnet in the overspecified case.
@@
Figure 5: Binomial simulation summary (response: LogLoss_Holdout versus true probabilities). Mean log-loss on a large holdout versus $n_{\text{total}}$, stratified by target $R^{2}$ and model order. All methods use non-relaxed paths (relax = FALSE) with the default mixing grid $\alpha\in\{0.5,1\}$. wSSE spikes at $n=p_{\mathrm{full}}$.
@@ 5 Conclusion
SVEMnet provides an open-source implementation of Self-Validated Ensemble Models for small-sample DOE and chemometric applications, with (relaxed) elastic-net base learners for Gaussian and binomial responses built on glmnet using validation-weighted information-criterion analogs.
Together with the permutation whole-model test and the random-search desirability scoring and candidate-selection tools, SVEMnet offers an end-to-end workflow for mixture-constrained multi-response optimization in a single toolchain. The software is aimed at practitioners running small-sample (e.g., tens of runs) chemometric, pharmaceutical, and industrial DOE studies for formulation optimization, often involving mixture factors alongside numeric and categorical process factors. SVEM-based elastic-net surrogates provide a practical alternative or complement to Gaussian-process-based Bayesian optimization, particularly when the number of runs is too small to support rich covariance structures or when full BO tooling is unavailable. A head-to-head comparison with Gaussian-process Bayesian optimization is beyond the scope of this software-focused paper; here our comparisons focus on repeated CV elastic-net baselines that reflect common penalized-regression workflows in small-sample DOE.

### v2

@@ Abstract
SVEMnet is an R package for fitting Self-Validated Ensemble Models (SVEM) with elastic-net base learners and performing multi-response optimization in small-sample mixture–process design-of-experiments (DOE) studies with numeric, categorical, and mixture factors. SVEMnet wraps elastic-net and relaxed elastic-net models for Gaussian and binomial responses from glmnet in a fractional random-weight (FRW) resampling scheme with anti-correlated train/validation weights; penalties are selected by validation-weighted AIC- and BIC-type criteria, and predictions are averaged across replicates to stabilize fits near the interpolation boundary. In addition to the core SVEM engine, the package provides deterministic high-order formula expansion, a permutation-based whole-model test heuristic, and a mixture-constrained random-search optimizer that combines Derringer–Suich desirability functions, bootstrap-based uncertainty summaries, and optional mean-level specification-limit probabilities to generate scored candidate tables and diverse exploitation and exploration medoids for sequential fit–score–run–refit workflows. A simulated lipid nanoparticle (LNP) formulation study illustrates these tools in a small-sample mixture–process DOE setting, and simulation experiments based on sparse quadratic response surfaces benchmark SVEMnet against repeated cross-validated elastic-net baselines.
@@ 1 Introduction
This paper introduces SVEMnet, an R package that implements SVEM with elastic-net and optional relaxed elastic-net base learners [8, 9, 10, 11]. In a relaxed fit, glmnet computes a penalized solution path indexed by the penalty parameter $\lambda$, identifies the active set at each $\lambda$, and then refits an unpenalized linear model on that active set, interpolating between penalized and unpenalized coefficients via a relaxation parameter $\gamma\in[0,1]$. In the sparse Gaussian linear model studied by Meinshausen [10], the relaxed lasso reduces the shrinkage of non-zero coefficients and can achieve lower prediction error than the standard lasso. Simulations in Section 4 indicate that combining relaxed base learners with the SVEM FRW ensemble can further reduce out-of-sample prediction error for Gaussian responses, plausibly because the ensemble averages over the additional variance introduced by relaxation.
@@
Self‑Validated Ensemble Models have already seen practical use in industrial DOE and chemometric settings, including mixture–process workflows for lipid nanoparticle (LNP) formulations and other pharmaceutical and analytical applications [17, 18, 6, 19, 20, 21, 22, 23, 24, 25]. These applications motivate an implementation tailored to chemometric DOE. In particular, SVEMnet is designed to automate many of the steps in the SVEM-based mixture–process workflow for LNP formulation optimization described by Karl et al. [17]. SVEMnet supplies an open-source SVEM implementation over glmnet with validation-weighted AIC/BIC selectors, deterministic high-order full model expansions, a permutation-based whole-model test heuristic, and a mixture-constrained random-search scoring tool for multi-response desirability and candidate generation.
@@ 2.2 SVEM background
following the FRW scheme for univariate analysis introduced by Xu et al. [3], which underpins the original SVEM extensions to penalized regression (and the introduction of the anti-correlated validation weights) by Gotwalt and Ramsey [4] and Lemkus et al. [2]. For each FRW replicate we then rescale both the training and validation weights to have mean one, $w_{i}\leftarrow w_{i}/\bar{w}$ for $w\in\{w^{\mathrm{train}},w^{\mathrm{valid}}\}$, matching glmnet’s internal normalization of observation weights.
Given $(w^{\mathrm{train}},w^{\mathrm{valid}})$ for replicate $b$, SVEMnet fits an elastic-net or relaxed elastic-net base learner to the FRW-weighted training sample under $w^{\mathrm{train}}$, evaluates a collection of path points using a validation-weighted information criterion computed with $w^{\mathrm{valid}}$ (Section 2.4), and stores predictions from the selected path point. Overall SVEM predictions are obtained by averaging these member predictions across $b=1,\dots,B$. Lemkus et al. [2] tuned penalties by validation-weighted SSE (wSSE); SVEMnet adopts the same FRW-and-ensemble structure but standardizes on validation-weighted AIC/BIC analogs—wAIC default for Gaussian responses; wBIC default for binomial responses.
@@
Related work by Lu and Anderson-Cook [33] adapts the SVEM FRW-based self-validation mechanism to define a self-validated mean squared prediction error (SVMSPE) for model comparison and pairs it with sequential non-uniform space-filling (NUSF) design augmentation that allocates more runs to regions where a small set of candidate models differ most in their predictions. In SVEMnet, the FRW validation losses instead feed validation-weighted AIC/BIC-type selectors within a given elastic-net SVEM model class, while sequential design is handled separately by the desirability-based random-search optimizer in Section 2.5.
@@ 2.4 Selection criterion
Because the FRW validation weights are random rather than fixed design weights, these scores are used heuristically for relative model comparison inside the FRW replicate rather than as exact AIC/BIC. As noted by Lu and Anderson-Cook [33], classical uses of AIC and BIC evaluate fit at the same observed locations used for estimation, which can lead to over-optimistic assessments. In SVEMnet, the wAIC and wBIC criteria are always computed on the FRW validation copy. For binomial models we retain the label wSSE even though the selector operates on the deviance scale rather than squared error.
@@ 2.5.2 Scoring and response weighting
User-specified goal weights $\tilde{w}_{r}$ are first normalized to $\bar{w}_{r}$ with $\sum_{r}\bar{w}_{r}=1$. When reweight_by_wmt = FALSE, we set $w_{r}=\bar{w}_{r}$ and report a single column score based on $S(x)$. When reweight_by_wmt = TRUE for Gaussian responses, we further reweight the normalized user weights by a monotone function of the permutation whole-model $p$-values from [5] (for example $\log_{10}(p_{r})$), renormalize to obtain WMT-adjusted weights $w_{r}^{\mathrm{WMT}}$, and compute a second geometric-mean score wmt_score by using $w_{r}=w_{r}^{\mathrm{WMT}}$ in $S(x)$. Methodologically, WMT compares the pattern of SVEM predictions on a space-filling set of design points with a reference distribution obtained by refitting the same SVEM specification to permuted responses, yielding an approximate whole-model $p$-value for the global null of a flat response surface [5]. These WMT-adjusted weights are diagnostic and do not alter the underlying SVEM fits. In this paper we treat the resulting WMT-based multipliers as a secondary diagnostic overlay: the primary optimization uses the user-specified weights $\tilde{w}_{r}$, and wmt_score is intended for sensitivity checks rather than as the sole ranking criterion.
@@ 5 Ionizable_Lipid_Type + N_P_ratio + flow_rate,
9 polynomial_order = 3, # up to cubic terms
@@ 3.1 Whole-model reweighting and permutation tests
To assess which responses show the strongest relationship with the study factors, we apply the permutation-based whole-model test heuristic (WMT) of Karl [5] to each SVEM model. WMT-based multipliers can optionally reweight the user-specified response weights, prioritizing responses with stronger predictive signal and down-weighting weaker ones. Consistent with Karl [5], we view these results as a way to highlight responses with clearer global signal, not as a hard filter: in small or noisy studies the WMT heuristic can have limited power, so candidate sets should not be chosen solely by appealing to WMT $p$-values or multipliers. Because WMT is permutation-based with finite simulation budgets, these $p$-values are approximate.
@@ 5 PDI = form_pdi),
Table 1: Permutation whole-model test results for the three responses (SVEMnet). Values are permutation-based and subject to Monte Carlo error.
@@
Size $1.00\text{\times}{10}^{-16}$
Potency $1.35\text{\times}{10}^{-14}$
PDI 0.72
@@ 3 Size = list(upper = 100),
The random-search scoring step then evaluates a large number of feasible settings (default 25,00025{,}000). For each candidate, svem_score_random() evaluates the SVEM models, computes per-response Derringer–Suich desirabilities, aggregates them into a single multi-response score via a weighted geometric mean, constructs a scalar uncertainty measure $U(x)$ from bootstrap percentile intervals, and, when specifications are supplied, computes mean-level probabilities that each candidate’s modeled mean lies inside the user limits. Supplying wmt = wmt_out adds a WMT-adjusted score column wmt_score in addition to the user-weighted score:
@@ 3.3 Computational environment and runtime
All timings were obtained on a Windows 11 x64 desktop with an Intel Core Ultra 7 265 CPU (20 cores, 2.4 GHz) and 64 GB RAM, running R 4.5.2 with SVEMnet 3.1.4 and glmnet 4.1-10 and parallel execution enabled for the WMT. Users should note that on Windows, the parallelization of the WMT will trigger a one-time Windows Firewall prompt for R when worker processes are created; this is due to local socket connections between R processes and does not involve external network access. The user may click Cancel on this alert window and the program will still run in parallel.
@@ 4.1 Design and metrics
Each replicate uses a sparse quadratic response surface in four continuous factors $X_{1}{:}X_{4}\in[-1,1]$ and one three-level factor $X_{5}\in{\mathrm{L1},\mathrm{L2},\mathrm{L3}}$ (sum-to-zero coding). Training data are Latin hypercube samples for $X_{1}{:}X_{4}$ with balanced $X_{5}$ levels [32], with $n_{\text{total}}\in\{15,20,\dots,50\}$. As noted in the introduction, we use Latin hypercubes here for simulation convenience; in practice, more sophisticated optimal or space-filling designs are typically preferable when constructing the initial experimental design [41, 28, 30, 31, 29].
@@
To reduce Monte Carlo error, the summary curves in Figures 3, 4, and 5 average over 1000 independent simulation replicates for each combination of $n_{\text{total}}$, target $R^{2}$, model order, and fitting method (Setting). In these overview plots we fix a common set of tuning defaults. For Gaussian responses, cv.glmnet uses standard (non-relaxed) elastic-net paths (relax = FALSE) and SVEMnet uses relaxed base learners (relax = TRUE); post-hoc debiasing is disabled for all methods and the elastic-net mixing grid is held at $\alpha\in\{0.5,1\}$. For binomial responses, all methods use non-relaxed paths (relax = FALSE) with the same default $\alpha$ grid. These settings are used for all model orders shown in the figures.
@@ 4.2 Gaussian results
Figure 3: Effect of expansion order on mean log-NRMSE (lower is better) plotted against $n_{\text{total}}$ for the Gaussian simulations. Rows correspond to target $R^{2}\in\{0.5,0.9\}$ and columns to the model order used in fitting. Colors and line types indicate the selector: repeated-cv.glmnet (blue dashed), wAIC (orange dotted), wBIC (green solid), and the loss-only wSSE selector (purple dot–dash). Light gray vertical lines in the middle and right-hand panels mark the interpolation boundaries $n_{\text{total}}=p_{\mathrm{full}}$ for Orders 2 and 3 ($p_{\mathrm{full}}=25$ and $p_{\mathrm{full}}=45$, respectively). Simulation settings match Figure 3: non-relaxed cv.glmnet, relaxed SVEMnet, no debiasing. The wSSE selector exhibits a pronounced spike in holdout log-NRMSE near $n_{\text{total}}=p_{\mathrm{full}}$, whereas wAIC and wBIC remain stable across the interpolation boundary.
Figure 4: Mean median-selected size $k_{\lambda}$ versus $n_{\text{total}}$ for the Gaussian simulations, stratified by target $R^{2}$ (rows) and model order (columns). Colors and line types indicate the selector: repeated-cv.glmnet (blue dashed), wAIC (orange dotted), wBIC (green solid), and the loss-only wSSE selector (purple dot–dash). The light gray vertical lines in the middle and right-hand panels mark the interpolation boundaries $n_{\text{total}}=p_{\mathrm{full}}$ for Orders 2 and 3 ($p_{\mathrm{full}}=25$ and $p_{\mathrm{full}}=45$, respectively). Simulation settings match Figure 3: non-relaxed cv.glmnet, relaxed SVEMnet, no debiasing. The wSSE selector exhibits a pronounced spike in $k_{\lambda}$ near $n_{\text{total}}=p_{\mathrm{full}}$.
@@ 4.3 Binomial results
Across $n_{\text{total}}\in\{15,\dots,50\}$ and $R^{2}\in\{0.5,0.9\}$, non-relaxed fits dominate their relaxed counterparts on the holdout log-loss scale. For the correctly specified quadratic expansion (Order 2), SVEMnet with wBIC and the non-relaxed repeated-cv.glmnet baselines perform similarly: Tukey–Kramer comparisons based on the mixed-effects model place these configurations in the same best-performing group (Table LABEL:tab:tukey-binomial). However, Figure 5 suggests that SVEMnet with wBIC may outperform cv.glmnet in the overspecified case.
@@
Figure 5: Binomial simulation summary (response LogLoss_Holdout versus true probabilities). Mean log-loss on a large holdout versus $n_{\text{total}}$, stratified by target $R^{2}$ (rows) and model order (columns). Colors and line types indicate the selector: repeated-cv.glmnet, wAIC, wBIC, and the loss-only wSSE selector. All methods use non-relaxed paths (relax = FALSE) with the default mixing grid $\alpha\in\{0.5,1\}$. Light gray vertical lines mark the interpolation boundaries $n_{\text{total}}=p_{\mathrm{full}}$ for Orders 2 and 3 ($p_{\mathrm{full}}=25$ and $p_{\mathrm{full}}=45$), where wSSE shows a pronounced spike in log-loss.
@@ 5 Conclusion
SVEMnet provides an open-source implementation of Self-Validated Ensemble Models for small-sample DOE and chemometric applications, with (relaxed) elastic-net base learners for Gaussian and binomial responses built on glmnet using validation-weighted information-criterion analogs. Together with the permutation whole-model test and the random-search desirability scoring and candidate-selection tools, SVEMnet offers an end-to-end workflow for mixture-constrained multi-response optimization in a single toolchain. The software is aimed at practitioners running small-sample (e.g., tens of runs) chemometric, pharmaceutical, and industrial DOE studies for formulation optimization, often involving mixture factors alongside numeric and categorical process factors.